# MoE-FFD: Mixture of Experts for Generalized and Parameter-Efficient Face Forgery Detection

### v2

@@ Abstract
Deepfakes have recently raised significant trust issues and security concerns among the public. Compared to CNN face forgery detectors, ViT-based methods take advantage of the expressivity of transformers, achieving superior detection performance. However, these approaches still exhibit the following limitations: (1) Fully fine-tuning ViT-based models from ImageNet weights demands substantial computational and storage resources; (2) ViT-based methods struggle to capture local forgery clues, leading to model bias; (3) These methods limit their scope on only one or few face forgery features, resulting in limited generalizability. To tackle these challenges, this work introduces Mixture-of-Experts modules for Face Forgery Detection (MoE-FFD), a generalized yet parameter-efficient ViT-based approach. MoE-FFD only updates lightweight Low-Rank Adaptation (LoRA) and Adapter layers while keeping the ViT backbone frozen, thereby achieving parameter-efficient training. Moreover, MoE-FFD leverages the expressivity of transformers and local priors of CNNs to simultaneously extract global and local forgery clues. Additionally, novel MoE modules are designed to scale the model’s capacity and smartly select optimal forgery experts, further enhancing forgery detection performance. Our proposed learning scheme can be seamlessly adapted to various transformer backbones in a plug-and-play manner. Extensive experimental results demonstrate that the proposed method achieves state-of-the-art face forgery detection performance with significantly reduced parameter overhead. The code is released at: https://github.com/LoveSiameseCat/MoE-FFD.
@@ 1 Introduction
Early traditional Deepfake detection methods focused on extracting hand-crafted features, such as eye-blinking frequency [2] and headpose inconsistency [3]. These techniques fail to capture representative features because the hand-crafted features may limit their scopes to only one or few kinds of statistical information. Heading into the era of deep learning, numerous CNN-based methods have been proposed to improve the detection accuracy [4, 5, 6, 7, 8]. Many of them employ Xception network [9] or EfficientNet [10] as backbones due to their outstanding performance in face forgery detection. To enhance generalizability and robustness, some methods proposed to extract common forgery featrues such as noise patterns [11, 12, 13], blending artifacts [4, 14], frequency fingerprints [15, 16, 17, 18], and identity inconsistency [8], etc. However, these methods are inherently limited to the local-interactions of CNN architectures.
@@
With the advent of the vision transformer (ViT) [19], ViT architectures have achieved significant success in a wide variety of computer vision tasks due to their long-range interactions and outstanding expressivity. In the realm of Deepfake detection, numerous ViT-based approaches [20, 21, 22, 23, 24, 25, 26] have been proposed, achieving enhanced accuracy and generalizability. For instance, Dong et al. [20] designed a ViT to capture the identity inconsistency between inner and outer face region. Zhuang et al. [22] proposed a ViT to mine unsupervised inconsistency-aware features within each frame. Shao et al. [21] introduced a concise yet effective Seq-DeepFake Transformer to predict a sequential vector of facial manipulation operations.
@@
Nevertheless, ViT-based forgery detection approaches still face several limitations. First, fully training ViT-based models from the ImageNet weights demands substantial computational resources, which hinders their deployment or fine-tuning in real-world applications, particularly on mobile devices with limited processing power. Second, although ViT-based methods exhibit outstanding expressivity, they may struggle to capture forgery features in local abnormal regions, resulting in model bias and limited generalizability. Finally, previous methods focus on certain forgery artifacts, but it is challenging to empirically select the optimal features in unpredictable application scenarios.
@@
This paper presents a generalized yet parameter-efficient approach MoE-FFD, which proposes using Mixture of Experts for Face Forgery Detection. MoE-FFD draws inspiration from Parameter Efficient Fine-Tuning (PEFT), which integrates lightweight Low-Rank Adaptation (LoRA) layers and Adapter layers with the ViT backbone. During the training process, only the designed LoRA and Adapter parameters are updated while the ViT parameters remain frozen as the ImageNet weights. The designed MoE modules dynamically select optimal LoRA and Adapter experts for face forgery detection.
@@
Compared with previous arts that directly fine-tune the entire ViT parameters, the proposed fine-tuning strategy effectively preserves abundant knowledge from ImageNet and enables the model to adaptively learn forgery-specific features. Additionally, the designed LoRA layers model the long-range interactions within input faces, while the Convpass Adapter layers effectively highlight local forgery anomalies. To this end, the integration of the designed LoRA and Adapter layers leverages the expressivity of transformers and the local forgery priors of CNNs, leading to enhanced generalizability and robustness. Additionally, we design novel Mixture of Experts (MoE) modules within both LoRA and Adapter layers to scale the model’s capacity using fixed activated parameters. The MoE dynamically selects optimal forgery detection experts for input faces, further enhancing the model’s performance.
@@
As depicted in Fig. 1, our MoE-FFD with the fewest activated parameters achieves the best AUC score on the unseen CelebDF-v2 datset [27]. Overall, the contributions of this work are summarized as follows:
@@
  • We innovatively integrate LoRA and Adapter modules with the ViT backbone for face forgery detection. This design enables the designed model to simultaneous mine global and local forgery clues in a parameter-efficient manner.
@@
  • We conduct experiments on six Deepfake datasets and various common perturbations. Our experimental results demonstrate that MoE-FFD achieves state-of-the-art generalizability and robustness. Extensive ablation experiments validate the effectiveness of our proposed MoE learning scheme and the designed components.
In this paper, Section 2 comprehensively reviews previous related literature. Section 3 details the proposed mixture of expert framework for face forgery detection. Section 4 presents extensive experimental results under various settings. Finally, Section 5 concludes the paper and discusses possible future research directions.
@@ 2.1 Face Forgery Detection
Early attempts at face forgery detection primarily relied on extracting handcrafted features, such as the lack of eye blinking [2], inconsistency of head pose [3], face warping artifacts [28], and heart rate anomalies [29]. However, these methods suffer from limited accuracy due to their narrow focus on specific statistical information. In response, learning-based approaches have emerged, leveraging generic network architectures like Xception Network [9], EfficientNet [10], and Capsule Network [30] for forgery feature extraction. Nonetheless, CNN-based methods are prone to overfitting to the training data, resulting in limited generalizability and robustness. Follow-up works, such as Face X-ray [4], F3Net [18], SBI [31], DCL [32], and RECCE [33], have introduced comprehensive forgery frameworks and robust feature extraction techniques, enhancing the model’s generalization capability. With the explosive development of ViT [19], numerous ViT-based methods [20, 21, 22, 23, 24, 25, 26] have been proposed to tackle the Deepfake problem. These approaches take advantage of non-inductive bias and global context understanding, achieving superior face forgery detection performance. However, most of these methods fine-tune the entire ViT model on Deepfake datasets, which is computationally expensive and may lead to loss of valuable knowledge from the ImageNet dataset [34]. In contrast, our approach finetunes external lightweight LoRA and Adapter parameters during training, keeping the ViT backbone fixed with ImageNet weights, thus enabling the model to learn forgery-specific knowledge and enhance the detection performance.
@@ 2.2 Parameter Efficient Fine-Tuning (PEFT)
Over the past few years, the size of deep learning models has exponentially increased, especially after the advent of transformers. Consequently, numerous Parameter Efficient Fine-Tuning (PEFT) methods have been proposed to reduce the computational and storage overhead. PEFT only updates a small portion of the model parameters while freezing the majority of pretrained weights. Adapter [35, 36] is a typical PEFT method, comprising a down-sample layer and an up-sample layer, generally integrated into transformer layers and blocks. Low-Rank Adaptation (LoRA) [37] layer aims at updating two low-rank matrices, significantly reducing the trainable parameters. Zhong et al. [38] found that convolution operation effectively introduces local prior information for image segmentation. Additionally, Visual Prompt tuning (VPT) [39] augments inputs with extra learnable tokens, which can be regarded as learnable pixels to vision transformers. Neural Prompt Search (NOAH) [40] advances further by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Scale and Shift Feature Modulation (SSF) [41] introduces to scale and shift parameters to modulate visual features during training, achieving comparable performance compared with full fine-tuning. Convpass [42] integrates convolutional bypasses into large ViT models, leveraging local priors to improve image classification performance. However, the application of PEFT methods to face forgery detection remains largely unexplored. In this work, we incorporate dedicated forgery Adapter and LoRA layers to mine global and local clues and achieve a parameter-efficient face forgery detection.
@@ 3.1.1 Vision Transformer
Vision Transformer consists of multiple blocks of Self-Attention and Multi-Layer Perceptron (MLP). For a given input sequence $x\in\mathbb{R}^{N_{t}\times D}$, $x$ is firstly projected to queries $Q\in\mathbb{R}^{N_{t}\times dim}$, keys $K\in\mathbb{R}^{N_{t}\times dim}$, and values $V\in\mathbb{R}^{N_{t}\times dim}$ using three learnable matrices $W_{q}\in\mathbb{R}^{D\times dim}$, $W_{k}\in\mathbb{R}^{D\times dim}$, and $W_{v}\in\mathbb{R}^{D\times dim}$, where $N_{t}$, $D$, and $dim$ denote the token number, embedding dimension, and hidden dimension, respectively. The $Q$, $K$, and $V$ are calculated by:
@@ 3.2 Overview of the Proposed Framework
Fig. 2 illustrates the designed MoE-FFD framework. Fig. 2 (a) depicts the ViT backbone, which is initialized with the ImageNet weights. The input face firstly undergoes the patch and position embedding. Then, the input tokens are delivered to the designed transformer blocks. Fig. 2 (b) presents more details regarding the designed transformer blocks. We modify the standard ViT blocks by integrating the designed MoE Adapter layer and MoE LoRA layer with the original blocks.
Fig. 2 (c) details the designed MoE Adapter layer. During the training process, the MLP parameters are frozen with the ImageNet weights. Within this layer, a MoE module is designed, comprising one gate $G_{A}(\cdot)$ and $M$ forgery Adapter experts. The gating mechanism aims to dynamically select the appropriate experts for each input query, while the designed Adapter experts aim to extract specific local forgery features from the input faces. Fig. 2 (d) illustrates the structure of Adapter $j$. We first reshape the input feature and reduce its channel dimension by using a 1$\times$1 convolution. Subsequently, we design convolution operations to extract specific local forgery clues. The output feature is then passed through to another 1$\times$1 convolution layer to restore the channel dimension. Finally, the feature map is flattened to the original shape.
Similarly, the attention weights $W_{q,k,v}$ in the MoE LoRA layer are fixed to the ImageNet weights. As shown in Fig. 2 (e), the MoE LoRA module consists of one gate $G_{L}(\cdot)$ and $N$ experts, each with a unique rank. The gate $G_{L}(\cdot)$ selectively activate sparse experts for detecting face forgeries. Fig. 2 (f) depicts the LoRA structure, where $\otimes$ indicates matrix multiplication. The output feature is derived by multiplying the input with two learnable low-rank matrices. The resulting $x_{out}$ is the element-wise summation of the fixed weights and the learned MoE LoRA weights, which are further processed by a self-attention mechanism.
@@ 3.3 Details of the MoE
LoRA modules with different ranks tend to project the input tokens into various feature spaces. However, in uncontrolled deployment environments, it is challenging to manually predefine an ideal rank for different testing faces. To tackle this challenge, we design a MoE LoRA Layer that learns an optimal LoRA expert for each input query. As shown in Fig. 2 (e), each LoRA expert $E_{L}(\cdot)_{i}$ specifies a rank $r_{i}$. In this work, we design a gating mechanism $G_{L}(\cdot)$ to dynamically select the Top-k experts (default: k=1). The details of the gating mechanism will be elaborated later. As such, the output tokens can be calculated by:
@@
TABLE I: Cross-dataset evaluation on five unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Methods highlighted in blue denote video-level results. #Params indicates the activated parameter number during training.
Method Venue #Params CDF WDF DFDC-P DFD DFR Avgerage
@@
Previous studies [21, 66, 67] have demonstrated the effectiveness of utilizing local difference convolution to capture face forgery clues. However, these CNN-based methods only limit their scopes in one specific forgery feature. In our research, we introduce the MoE Adapter Layer integrated into the ViT backbone. This design aims to scale the model capacity and facilitate the dynamic selection of the suitable forgery Adapter expert. Additionally, the designed MoE Adapter layer dynamically injects local forgery priors into the plain ViT backbone.
@@
In ADC and CDC, we set the $\Omega$ size as 3$\times$3. The $\hat{x}_{p}$ in CDC can be calculated by: $\hat{x}_{p}=x_{p}-x_{c}$. And the $\hat{x}_{p}$ in ADC is denoted as: $\hat{x}_{p}={x}_{p}-{x}_{p}^{next}$, where ${x}_{p}^{next}$ is the next element in the clockwise direction.
In the RDC and SOC operations, we set the size of the $\Omega$ region as 5$\times$5. For each element ${x}_{p}$, we define a corresponding radial element ${x}_{p}^{R}$ in the peripheral region of $\Omega$, which is highlighted in dark green in Fig. 3 (d) and (e). As such, the $\hat{x}_{p}$ in RDC is calculated by: $\hat{x}_{p}={x}_{p}^{R}-{x}_{p}$. SOC aims at learning second-order local anomaly of the input. As such, $\hat{x}_{p}$ in SOC is formulated as: $\hat{x}_{p}=(x_{p}^{R}-x_{p})-(x_{p}-x_{c})=(x_{p}^{R}-x_{p})+(x_{c}-x_{p})$. As such, the designed MoE Adapter layer can effectively search intrinsic detailed local forgery patterns in a larger feature space.
@@
We adopt Top-k noisy gating [48] as our gating mechanism. The gating scores $G(x)\in\mathbb{R}^{N_{e}}$ are determined by the values $H(x)\in\mathbb{R}^{N_{e}}$, where $N_{e}$ indicates the expert number. For a given input $x\in\mathbb{R}^{N_{t}\times dim}$, we first apply average pooling and reshape it to $x_{m}\in\mathbb{R}^{dim}$. Then, we calculate $H(x)$ by:
@@
where $W_{gate}\in\mathbb{R}^{dim\times N_{e}}$ and $W_{noise}\in\mathbb{R}^{dim\times N_{e}}$ are slim trainable parameters. $\rm Softplus$ is an activation function. For a given $H(x)\in\mathbb{R}^{N_{e}}$, we keep Top-k ($k\leq N_{e}$) values while setting others as $\infty$- ∞. Then, the gating scores $G(x)$ can be calculated by:
@@
$CV(Importance(X))=\frac{{\rm Mean}[Importance(X)]}{{\rm Std}[Importance(X)]}.$ (15)
As such, the additional loss $L_{moe}$ encourages all experts to have equal importance.

### v3

@@ Abstract
Deepfakes have recently raised significant trust issues and security concerns among the public. Compared to CNN-based face forgery detectors, ViT-based methods take advantage of the expressivity of transformers, achieving superior detection performance. However, these approaches still exhibit the following limitations: (1) Fully fine-tuning ViT-based models from ImageNet weights demands substantial computational and storage resources; (2) ViT-based methods struggle to capture local forgery clues, leading to model bias; (3) These methods limit their scope on only one or few face forgery features, resulting in limited generalizability. To tackle these challenges, this work introduces Mixture-of-Experts modules for Face Forgery Detection (MoE-FFD), a generalized yet parameter-efficient ViT-based approach. MoE-FFD only updates lightweight Low-Rank Adaptation (LoRA) and Adapter layers while keeping the ViT backbone frozen, thereby achieving parameter-efficient training. Moreover, MoE-FFD leverages the expressivity of transformers and local priors of CNNs to simultaneously extract global and local forgery clues. Additionally, novel MoE modules are designed to scale the model’s capacity and smartly select optimal forgery experts, further enhancing forgery detection performance. Our proposed learning scheme can be seamlessly adapted to various transformer backbones in a plug-and-play manner. Extensive experimental results demonstrate that the proposed method achieves state-of-the-art face forgery detection performance with significantly reduced parameter overhead in cross-dataset, cross-manipulation, and robustness evaluations. Our ablation studies further validate the effectiveness of the designed components and the proposed learning scheme. The code is available at: https://github.com/LoveSiameseCat/MoE-FFD.
@@ 1 Introduction
Early traditional Deepfake detection methods focused on extracting hand-crafted features, such as eye-blinking frequency [2] and headpose inconsistency [3]. These techniques fail to capture representative features because the hand-crafted features may limit their scopes to only one or few kinds of statistical information. Heading into the era of deep learning, numerous CNN-based methods have been proposed to improve the detection accuracy [4, 5, 6, 7, 8]. Many of them employ Xception network [9] or EfficientNet [10] as backbones due to their outstanding performance in face forgery detection. To enhance generalizability and robustness, some methods proposed to extract common forgery features such as noise patterns [11, 12, 13], blending artifacts [4, 14], frequency fingerprints [15, 16, 17, 18], and identity inconsistency [8], etc. However, these methods are inherently limited to the local-interactions of CNN architectures.
@@
With the advent of the vision transformer (ViT) [19], ViT architectures have achieved significant success in a wide variety of computer vision tasks due to their long-range interactions and outstanding expressivity. In the realm of Deepfake detection, numerous ViT-based approaches [20, 21, 22, 23, 24, 25, 26] have been proposed, achieving enhanced accuracy and generalizability. For instance, Dong et al. [20] designed a ViT to capture the identity inconsistency between inner and outer face region. Zhuang et al. [22] proposed a ViT-based model to mine unsupervised inconsistency-aware features within each frame. Shao et al. [21] introduced a concise yet effective Seq-DeepFake Transformer to predict a sequential vector of facial manipulation operations.
@@
Nevertheless, ViT-based forgery detection approaches still face several limitations. First, fully training ViT-based models from the ImageNet weights demands substantial computational resources, which hinders their deployment or fine-tuning in real-world applications, especially on resource-constrained mobile devices. Second, although ViT-based methods exhibit outstanding expressivity, they may struggle to capture forgery features in local abnormal regions, resulting in model bias and limited generalizability. Finally, previous methods focus on certain forgery artifacts, but it is challenging to empirically select the optimal features in unpredictable application scenarios.
@@
This paper presents a generalized yet parameter-efficient approach MoE-FFD, which proposes using Mixture of Experts for Face Forgery Detection. MoE-FFD draws inspiration from Parameter Efficient Fine-Tuning (PEFT), which integrates lightweight Low-Rank Adaptation (LoRA) layers and Adapter layers with the ViT backbone. During the training process, only the designed LoRA and Adapter parameters are updated while the ViT parameters remain frozen at their initial ImageNet weights. The designed MoE modules dynamically select optimal LoRA and Adapter experts for face forgery detection.
@@
Compared with previous arts that directly fine-tune the entire ViT parameters, the proposed fine-tuning strategy effectively preserves abundant knowledge from ImageNet and enables the model to adaptively learn forgery-specific features. Additionally, the designed LoRA layers model the long-range interactions within input faces, while the Convpass Adapter layers effectively highlight local forgery anomalies. To this end, the integration of the designed LoRA and Adapter layers leverages the expressivity of transformers and the local forgery priors of CNNs, leading to enhanced generalizability and robustness. Additionally, we design novel Mixture of Experts (MoE) modules within both LoRA and Adapter layers, enabling the model to scale its capacity efficiently by activating only a subset of parameters dynamically. The MoE modules adaptively select optimal forgery detection experts based on input faces, further improving detection performance.
@@
As depicted in Fig. 1, our MoE-FFD with the fewest activated parameters achieves the best AUC score on the unseen CelebDF-v2 dataset [27]. Overall, the contributions of this work are summarized as follows:
@@
  • We innovatively integrate LoRA and Adapter modules with the ViT backbone for face forgery detection. This design enables the model to simultaneously mine global and local forgery clues in a parameter-efficient manner.
@@
  • We conduct experiments on seven Deepfake datasets and various common perturbations. Our experimental results demonstrate that MoE-FFD achieves state-of-the-art generalizability and robustness. Extensive ablation experiments validate the effectiveness of our proposed MoE learning scheme and the designed components.
The remainder of this paper is structured as follows. Section 2 comprehensively reviews previous related literature. Section 3 details the proposed mixture of expert framework for face forgery detection. Section 4 presents extensive experimental results under various settings. Finally, Section 5 concludes the paper and discusses potential directions for future research.
@@ 2.1 Face Forgery Detection
Early attempts at face forgery detection primarily relied on extracting handcrafted features, such as the lack of eye blinking [2], inconsistency of head pose [3], face warping artifacts [28], and heart rate anomalies [29]. However, these methods suffer from limited accuracy due to their narrow focus on specific statistical information. In response, learning-based approaches have emerged, leveraging generic network architectures like Xception Network [9], EfficientNet [10], and Capsule Network [30] for forgery feature extraction. Nonetheless, CNN-based methods are prone to overfitting to the training data, resulting in limited generalizability and robustness [31, 32]. Follow-up works, such as Face X-ray [4], F3Net [18], SBI [33], DCL [34], RECCE [35], and MVIM [36], have introduced comprehensive forgery frameworks and robust feature extraction techniques, enhancing the model’s generalization capability. ADA-FInfer [37] and Diff-ID [38] take one step forward to achieve explainable Deepfake detection. Recent LVLM-based works [39, 40, 41, 42, 43] aim to achieve explainable face forgery detection. With the explosive development of ViT [19], numerous ViT-based methods [20, 21, 22, 23, 24, 25, 26, 42] have been proposed to tackle the Deepfake problem. These approaches take advantage of non-inductive bias and global context understanding, achieving superior face forgery detection performance. However, most of these methods fine-tune the entire ViT model on Deepfake datasets, which is computationally expensive and may lead to loss of valuable knowledge from the ImageNet dataset [44]. In contrast, our approach finetunes external lightweight LoRA and Adapter parameters during training, keeping the ViT backbone fixed with ImageNet weights, thus enabling the model to learn forgery-specific knowledge and enhance the detection performance.
@@ 2.2 Parameter Efficient Fine-Tuning (PEFT)
Over the past few years, deep learning models have increased exponentially in size, especially after the advent of transformers. Consequently, numerous Parameter Efficient Fine-Tuning (PEFT) methods have been proposed to reduce the computational and storage overhead. PEFT only updates a small subset of the model parameters while freezing the majority of pretrained weights. Adapter [45, 46] is a typical PEFT method, comprising a down-sample layer and an up-sample layer, usually integrated into transformer layers and blocks. Low-Rank Adaptation (LoRA) [47] layer aims at updating two low-rank matrices, significantly reducing the trainable parameters. Zhong et al. [48] found that convolution operation effectively introduces local prior beneficial for image segmentation. Additionally, Visual Prompt Tuning (VPT) [49] augments inputs with extra learnable tokens, which can be regarded as learnable pixels to vision transformers. Neural Prompt Search (NOAH) [50] advances further by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Scale and Shift Feature Modulation (SSF) [51] introduces to scale and shift parameters to modulate visual features during training, achieving comparable performance compared with full fine-tuning. Convpass [52] integrates convolutional bypasses into large ViT models, leveraging local priors to improve image classification performance. However, the application of PEFT methods to face forgery detection remains largely unexplored. In this work, we incorporate dedicated forgery Adapter and LoRA layers to mine global and local clues, thereby achieving a parameter-efficient face forgery detection.
@@ 3.1.1 Vision Transformer
Vision Transformer consists of multiple blocks of self-attention and Multi-Layer Perceptron (MLP). For a given input sequence $\in\mathbb{R}^{N_{t}\times D}, xx$ is firstly projected to queries $Q\in\mathbb{R}^{N_{t}\times dim}$, keys $K\in\mathbb{R}^{N_{t}\times dim}$, and values $V\in\mathbb{R}^{N_{t}\times dim}$ using three learnable matrices $W_{q}\in\mathbb{R}^{D\times dim}$, $W_{k}\in\mathbb{R}^{D\times dim}$, and $W_{v}\in\mathbb{R}^{D\times dim}$, where $N_{t}$, $D$, and dimdim denote the token number, embedding dimension, and hidden dimension, respectively. The $Q$, $K$, and $V$ are calculated by:
@@ 3.2 Overview of the Proposed Framework
Fig. 2 illustrates the designed MoE-FFD framework. Fig. 2 (a) depicts the ViT backbone, which is initialized with the ImageNet weights. The input face is first processed through patch and positional embedding, after which the embedded tokens are passed into the customized transformer blocks. Fig. 2 (b) presents more details regarding the designed transformer blocks. We modify the standard ViT blocks by integrating the designed MoE Adapter layer and MoE LoRA layer with the original blocks.
Fig. 2 (c) details the designed MoE Adapter layer. During the training process, the MLP parameters are frozen with the ImageNet weights. Within this layer, an MoE module is designed, comprising one gate $G_{A}(\cdot)$ and $M$ forgery Adapter experts. The gating mechanism aims to dynamically select the appropriate experts for each input query, while the designed Adapter experts aim to extract specific local forgery features from the input faces. Fig. 2 (d) illustrates the structure of Adapter $j$. We first reshape the input feature and reduce its channel dimension by using a $\times1$ convolution. Subsequently, we design convolution operations to extract specific local forgery clues. The output feature is then passed through to another $\times1$ convolution layer to restore the channel dimension. Finally, the feature map is flattened to the original shape.
Similarly, the attention weights $W_{q,k,v}$ in the MoE LoRA layer are fixed to the ImageNet weights. As shown in Fig. 2 (e), the MoE LoRA module consists of one gate $G_{L}(\cdot)$ and $N$ experts, each with a unique rank. The gate $G_{L}(\cdot)$ selectively activate sparse experts for detecting face forgeries. Fig. 2 (f) depicts the LoRA structure, where $\otimes$ indicates matrix multiplication. The output feature is derived by multiplying the input with two learnable low-rank matrices. The resulting $x_{out}$ is the element-wise summation of the fixed weights and the learned MoE LoRA weights, which are further processed by a self-attention mechanism. Overall, the MoE process can be summarized as follows: 1) The input feature $x_{in}$ is passed into the MoE module; 2) A gating network processes the input and outputs a set of logits, one for each expert; 3) Top-k (k=1 in this work) gating selects the most relevant expert; 4) The input is then routed to the selected expert and output $x_{out}$.
@@ 3.3 Details of the MoE
LoRA modules with different ranks tend to project the input tokens into various feature spaces. However, in uncontrolled deployment environments, it is challenging to manually predefine an ideal rank for different testing faces. To tackle this challenge, we design an MoE LoRA Layer that learns an optimal LoRA expert for each input query. As shown in Fig. 2 (e), each LoRA expert $E_{L}(\cdot)_{i}$ specifies a rank $r_{i}$. In this work, we design a gating mechanism $G_{L}(\cdot)$ to dynamically select the Top-k experts (default: k=1). The details of the gating mechanism will be elaborated later. As such, the output tokens can be calculated by:
@@
TABLE I: Cross-dataset evaluation on five unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Methods in the bottom table denote video-level results. #Params indicates the activated parameter number during training.
Method Venue #Params CDF WDF DFDC-P DFD DFR AVG
@@
Previous studies [21, 79, 80, 81, 82] have demonstrated the effectiveness of utilizing local difference convolution to capture face forgery clues. However, these CNN-based methods only limit their scopes in one specific forgery feature. In our research, we introduce the MoE Adapter Layer integrated into the ViT backbone. This design aims to scale the model capacity and facilitate the dynamic selection of the suitable forgery Adapter expert. Additionally, the designed MoE Adapter layer dynamically injects local forgery priors into the plain ViT backbone.
@@
In ADC and CDC, we set the $\Omega$ size as $\times3$. The $\hat{x}_{p}$ in CDC can be calculated by: $\hat{x}_{p}=x_{p}-x_{c}$. And the $\hat{x}_{p}$ in ADC is denoted as: $\hat{x}_{p}={x}_{p}-{x}_{p}^{next}$, where ${x}_{p}^{next}$ is the next element in the clockwise direction.
In the RDC and SOC operations, we set the size of the $\Omega$ region as $\times5$. For each element ${x}_{p}$, we define a corresponding radial element ${x}_{p}^{R}$ in the peripheral region of $\Omega$, which is highlighted in dark green in Fig. 3 (d) and (e). As such, the $\hat{x}_{p}$ in RDC is calculated by: $\hat{x}_{p}={x}_{p}^{R}-{x}_{p}$. SOC aims at learning second-order local anomaly of the input. As such, $\hat{x}_{p}$ in SOC is formulated as: $\hat{x}_{p}=(x_{p}^{R}-x_{p})-(x_{p}-x_{c})=(x_{p}^{R}-x_{p})+(x_{c}-x_{p})$. As such, the designed MoE Adapter layer can effectively search intrinsic detailed local forgery patterns in a larger feature space.
@@
The gating network in a Mixture-of-Experts (MoE) model dynamically routes each input to the most appropriate expert, determining the contribution of each expert to the final prediction. It performs dynamic routing, enabling the MoE model to adaptively select appropriate experts that are most relevant for each input sample. We adopt Top-k noisy gating [58] as our gating mechanism. The gating scores $G(x)\in\mathbb{R}^{N_{e}}$ are determined by the values $H(x)\in\mathbb{R}^{N_{e}}$, where $N_{e}$ indicates the expert number. For a given input $x\in\mathbb{R}^{N_{t}\times dim}$, we first apply average pooling and reshape it to $x_{m}\in\mathbb{R}^{dim}$. Then, we calculate $H(x)$ by:
@@
where $W_{gate}\in\mathbb{R}^{dim\times N_{e}}$ and $W_{noise}\in\mathbb{R}^{dim\times N_{e}}$ are slim trainable parameters. $\rm Softplus$ is an activation function. For a given $H(x)\in\mathbb{R}^{N_{e}}$, we keep Top-k ($k\leq N_{e}$) values while setting others as -$\infty$. Then, the gating scores $G(x)$ can be calculated by:
@@
$CV(Importance(X))=\frac{{\rm Std}[Importance(X)]}{{\rm Mean}[Importance(X)]}.$ (15)
As such, the additional loss $L_{moe}$ encourages all experts to have equal importance and allows the model to avoid relying on a single expert during training, which can effectively prevent overfitting in face forgery detection.